OpenStudy (anonymous):

Express as a single logarithm 2lnx-1/2 ln (x^2-1) + 3 ln (x^2 +1)

OpenStudy (mertsj):

\[\frac{2\ln (x-1)}{2\ln (x^2-1)}+3\ln (x^2+1)\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

no , it is 2lnx ) - (1/2) (x^2-1) + 3ln (x^2+1)

OpenStudy (anonymous):

\[2lnx - \frac{1}{2}\ln(x^2-1) + 3\ln(x^2 + 1)\]Is this it?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

\[\ln x^2-\ln (x^2-1)^{\frac{1}{2}}+^{}\ln (x^2+1)^3=\ln x^2+\ln (x^2+1)^3-\ln (x^2-1)^{\frac{1}{2}}\]

OpenStudy (mertsj):

\[\ln x^2(x^2+1)^3-\ln (x^2-1)^{\frac{1}{2}}=\ln \frac{x^2(x^2+1)^3}{(x^2-1)^{\frac{1}{2}}}\]

OpenStudy (anonymous):

i get it , thanks